Mr odongo is an inventory manager for office supplies in a large office supply warehouse.

The annual demand for paper punches is 20000 units. The ordering cost is $100 per order and carrying(holding) cost is $5 per unit per year. Determine the expected ordering quantity and the fixed order


1
Expert's answer
2021-09-27T15:05:12-0400

Solution:

Given: Annual demand=20000

Order quantity=100

Carrying cost=5

Expected ordering quantity"=\sqrt{\dfrac{2\times Annual\ demand \times order\ quantity}{carrying\ cost}}"

"=\sqrt{\dfrac{2\times 20000\times 100}{5}}\approx 894"

The fixed order cannot be calculated as data is insufficient.